Next Section | Table of Contents


Site 1204
Transit to Site 1204 and Underway Survey

The provisional location of Site 1204 was at or close to Site 883, where ˜38 m of basalt had been penetrated and partly cored during Leg 145. The site survey was under way 1 hr after departing Site 1203.

After the vessel slowed to 6 kt, seismic equipment and a magnetometer were deployed and the survey started using one 80-in3 water gun and a single-channel Teledyne hydrophone streamer. Good-quality reflection profiles were displayed on two software-controlled EPC recorders operating at different scanning speeds. The position selected was ˜0.5 km from the six holes cored and drilled during Leg 145 at Site 883. At 1500 hr on 25 July 2001, the seismic equipment was retrieved as the vessel came about to return to Site 1204. Thrusters were lowered and the vessel maneuvered to the Global Positioning System (GPS) coordinates of the new location by 1534 hr. The hydrophones were then lowered concurrent with the making up of a new BHA. A beacon was deployed at 1614 hr. The 37-nmi survey was accomplished in 6.2 hr at an average speed of 6.0 kt.

Hole 1204A

After the vessel settled on location, the corrected PDR depth referenced to the DES was obtained and indicated 2604.4 m. The BHA was made up of a 9.875-in RCB, a medium-hard formation C-4 bit, a mechanical bit release, a modified head sub, an outer core barrel, a modified top sub, a modified head sub, seven 8.25-in drill collars, a tapered drill collar, six 5.5-in drill pipe sections, and one crossover sub.

Hole 1204A was spudded with the RCB at 2015 hr on 25 July. The drill string was drilled ahead with a center bit in place to a depth of 761.9 mbsf, where coring was initiated in nannofossil ooze and chalk. Coring advanced through 57.6 m of sediment at an average ROP of 28.8 m/hr with 66.3% recovery. Basement coring began with Core 197-1204A-7R (819.5–829.0 mbsf). As coring progressed, a low-pressure system developed and the weather deteriorated. By the early morning hours of 28 July, vessel heave was occasionally measured at 16 ft. Winds were 25–30 kt from the northwest, and seas were estimated between 10 and 12 ft and from the northwest. The main swell emanated from the north-northeast. The barometer was steady at 1000 Mb.

After Core 197-1204A-10R (843.3–857.9 mbsf) (ROP = 2.0 m/hr) was recovered with 79% recovery (fresh basalt), seawater was observed flowing back at the rig floor connection, indicating that the float valve was jammed open (probably with basalt fragments). Three runs were made with the deplugger before the flow was stopped. Coring resumed with Core 197-1204A-11R (857.9–862.7 mbsf), representing a 4.8-m advance. No recovery was achieved, but the appearance of the core catcher with one missing finger suggested that the core had possibly slipped out during the recovery of the core barrel. A fresh core barrel was dropped, and Core 197-1204A-12R (862.7–867.5 mbsf) (ROP = 3.8 m/hr) was cored an additional 4.8 m. The core barrel was recovered empty. A deplugger was deployed with paint on the latch dogs to verify the engagement of the core barrel within the BHA. The deplugger was recovered and appeared to have engaged in the latch sleeve. It was surmised that a formation change signaled by the increased ROP could have involved the coring of volcaniclastic sediment, and because of the large heave, the sediment was not recovered. A fresh core barrel was dropped and Core 197-1204A-13R (867.5–877.1 mbsf) (ROP = 2.6 m/hr) advanced 9.6 m. The core barrel was recovered empty.

In a last attempt to save the hole, a fifth deplugger run was made and engagement with the latch sleeve was suggested by the abrasion on the latch dogs of the rotary assembly. This time, Core 197-1204A-14R (877.1–880.3 mbsf) (ROP = 5.5 m/hr) was intentionally advanced only 3.2 m. When there was no recovery, the drill string was tripped to the surface to determine the mechanical reason for the failure to recover core. Because of rough sea conditions, the deployment of a FFF was not attempted. Total core recovered from Hole 1204A was 55.86 m and recovery was 47.2%.

When the BHA was recovered, the bit was removed and the throat of the bit was found to be full of basaltic debris, some of which were current-rounded pebbles. The bottom of the mechanical bit release below the float valve was also jammed with basalt. The total amount of recovered basalt filled one 1.5-m section. The bit was found to be in very good condition (1, 1, NO, E, 1/16, HP, RR) after accumulating 30.3 total rotating hours. Of the total rotating hours, 10 hr were required to drill 762 m and an additional 2 hr needed to core the ˜58 m of sediment. The remaining 18.3 hr was expended penetrating 60 m into basement.

The latch sleeve was inspected and found to be in gage with no apparent wear. A new C-4 bit was affixed to a fresh mechanical bit release and the drill string deployed to spud Hole 1204B. This incident was classified as weather induced. The AHC was in operation during this incident.

Hole 1204B

The vessel was offset 100 m southeast of Hole 1204A, and Hole 1204B was spudded at 2015 hr on 28 July. The seafloor was tagged at 2381.0 m relative to the DES on the rig floor.

The RCB system was drilled ahead with an average ROP of 73.7 m/hr to a depth of 810.7 mbsf. This depth was estimated to be within 5–10 m of basement. The first core (197-1204B-1R [810.7–820.3 mbsf]) contacted basement at a depth of 816.0 mbsf. Coring then advanced without incident to a depth of 906.3 mbsf. After recovering Core 197-1204B-11R, seawater was flowing back at the rig floor level from the open end of the drill string, indicating that basalt had jammed in the float valve. This was the same phenomenon that forced the early cancellation of Hole 1204A. It required six deployments of the deplugger and one core barrel round trip to clear the obstruction from the float valve. The time expended in this process was 8.8 hr.

Coring resumed, and Hole 1204B was deepened to 954.5 mbsf when time expired on site. A total of 138.5 m of basaltic basement was cored with an average recovery of 38.0%. The cored sedimentary section was 5.3 m with 62.5% recovery. The average ROP in basement was 3.6 m/hr. The basement ROP ranged from 0.9 m/hr (Core 197-1204B-9R [887.1–888.8 mbsf]) to 5.5 m/hr in the bottom 19.3 m of the hole. An operation problem required using the aft coring line to retrieve Core 197-1204B-17R. Total core recovered from Hole 1204B was 55.93 m (recovery = 38.9%).

The drilling equipment was secured in preparation for the 2-day transit to Site 1205. This included the complete disassembly of the BHA. Although the bit had acquired 50.4 hr, it appeared to be in excellent shape with only two plugged nozzles to show that it had been in service. The bit was slightly under gage with all the inserts in the cones and had no chipped teeth. The beacon was successfully recovered. The vessel began the transit to the next site at 2100 hr on 1 August.

Next Section | Table of Contents